# Insights and Current Gaps in Open-Source LLM Vulnerability Scanners: A Comparative Analysis

### v1

@@ IV-F Summary of Distinctive Features and their Comparison
AI Scanner Name Test-suite Focus Guardrails Interface Explainability Capabilities Multi-language Support Customizable Attacks Insecure Coding Tests
Garak String-matching $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$
PyRIT LLM-based $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$
Giskard LLM-based $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
CyberSec. Pattern-matching $\times$ $\times$ $\times$ $\times$ $\checkmark$
@@ V-A Evaluation Objectives
This evaluation provides a comparative analysis of the scanners described in the previous section, focusing on two key aspects: attack coverage and evaluator effectiveness.
Attack Coverage. A scanner’s comprehensiveness is determined by its attack coverage, which is essential for effectively simulating a wide range of real-world threats. We evaluate attack coverage by examining the diversity of attack types; the volume of attack instances produced for each type; and the quality of these attacks, measured by their effectiveness in triggering errors in the target LLM.
Evaluator Effectiveness. The reliability of a scanner can be assessed by the effectiveness of its evaluators. We assesses evaluators’ accuracy in detecting vulnerabilities for various attack scenarios.
@@ V-B Evaluation Methodology and Results
To assess the scanners’ effectiveness in identifying vulnerabilities, each scanner was tested against four LLMs: Meta’s LLaMA 3 [1], Cohere’s Command-R [3], OpenAI’s GPT-4o [2], and Mistral AI’s Mistral Small [52]. To assess the quality of the attacks provided by each scanner, we applied a comprehensive series of these attacks and measured their success rates (ASR) using each scanner evaluator. However, evaluators can also make errors, necessitating the assessment of their margin of error (MOE) to ensure reliable attack success rates reported. Due to the lack of inherent ground truth in scanner evaluations, we manually annotated over a thousand attack responses to establish a reliable baseline. This manual annotation process required a thorough understanding of each attack’s objective, determining its success by analyzing the model’s response. We ensured balanced representation across attack types, facilitating the calculation of the evaluators’ accuracy and their MOE. Further details regarding our evaluation and data curation methodology are in the Appendix C.
TABLE II: Attacks’ success rate (ASR) and reliability (MOE) over different LLM models.
@@
Table II shows for each scanner its different attack categories, alongside their number of prompt instances, success rates (ASR) and margin of error (MOE) across the four tested LLMs. Figure 4 shows the ASR and MOE by the attack perspective. For the sake of comparability, the scanners’ partially overlapping $\sim$35 attack types where grouped into five categories (Jailbreak, Gradient-based, Context and Continuation C&C, Insecure Code Generation CG and General Multi-turn) - see details in Appendix A. Garak stands out for providing the most extensive variety of attacks and, in most cases, the highest number of instances per attack. Additionally, Garak’s attacks are of the highest quality, with success rates ranging from approximately 20% in context and continuation attacks to nearly 70% in insecure code attacks. The MOE measurements reveal notable errors across all scanner evaluators, with PyRIT’s LLM-based evaluators being the most reliable (lowest MOE), while Garak shows a maximum MOE of 26%, corresponding to a 37% error rate in detecting successful attacks.
Both static and LLM-based evaluators experience this issue. Static evaluators that use string-matching can oversimplify, e.g., we found misclassifications due to minor punctuation changes - see Appendix C-B. LLM-based evaluators, while better at handling complex contexts, can be unpredictable, sometimes offering irrelevant reasoning. In one example (Appendix C-B), the evaluator LLM tests the presence of a specific safety mechanism instead of recognizing that the target LLM successfully evades the attack altogether, regardless of whether such a mechanism exists. As evident in Fig. 1, the LLM-based evaluators of Giskard and PyRIT offer the highest reliability. We believe CyberSecEval’s low reliability comes from the ambitious attempt at identifying unsafe code. As for Garak - we believe its large-scale data will benefit transitioning from static to LLM-based evaluators. Last but not least, Fig. 1 shows a clear advantage for static-focused scanners in terms of attack effectiveness.
@@
Figure 4: Per-attack performance, averaged over the target LLMs. MOE shown as error bars.
VI Recommendations
VI-A Matching Red-Teaming Scanners to Your Organizational Needs
Based on our review and scanner performance evaluation, this section offers strategic recommendations for selecting the most suitable scanner for specific organizational needs. Selecting the right scanner for organizational red-teaming efforts is a complex task, influenced by specific business risks.
@@ VI-A Matching Red-Teaming Scanners to Your Organizational Needs
Red-teaming groups handling diverse use cases, such as large corporations employing LLMs for HR, sales, marketing, internal data analysis, etc. would prefer a ready-to-use test suite offering wide coverage of potential vulnerabilities, where customization is less frequent. In contrast, agile firms focusing on single-case products would favor customizable do-it-yourself solutions that allow for dynamic adaptations through tailored tests that meet their specific needs.
Below we position the analyzed scanners on the ”ready-to-use” versus ”do-it-yourself” spectrum - highlighting their strengths and limitations to guide organizations in making informed choices.
@@
Garak offers the most extensive test-suite, making it suitable for red-teaming groups that deal with diverse use-cases. However, its focuses more on a static attack dataset, limiting customizability. Garak also integrates with Nvidia’s NeMo Guardrails, enabling setting additional safety layers.
@@
Giskard is ideal for users seeking flexible attack generation with both static and LLM-based methods. It offers a simple yet effective customization of LLM-based attacks. Hence, it enables producing tailored test suites for various attack types, ideal for red-teaming in dynamic online environments with minimal manual interference. Moreover, it includes Guardrails to enhance safety.
@@
PyRIT offers the most customizable test suite, focusing on LLM-based attacks. It allows users to edit both attacker and evaluator LLMs, providing full access to their instructions. This offers extensive flexibility but requires significant prompt engineering. Therefore, PyRIT is best suited for red-teams focusing on an internally crafted test-suite rather than relying on external knowledge. Additionally, PyRIT implements advanced multi-step attacks which exhibits relatively high success rate.
@@
CyberSecEval focuses on red-teaming for code-generating LLMs. Its test-suite is designed to expose code-related security issues. This makes it valuable for red-teaming groups dealing with generative AI for software and cybersecurity, where the integrity of auto-generated code is crucial.
VI-B Current Scanner Gaps and Suggested Improvements
We have discovered critical issues concerning the scanners’ reliability, stemming from the evaluator component’s performance. This component is typically tailored for the attack at hand, rendering existing datasets obsolete for scanner’s distinctive tests. To this end, we curated a small scale dataset of $\sim$1K samples designed to assess the correct classification of attacks’ success in the evaluated scanners. This revealed that all evaluators employed by the scanners were insufficiently robust, facing challenges in both static and LLM-based evaluators. To improve the reliability of evaluators and, by extension, the scanners themselves, we propose several key initiatives: Develop a Benchmarking Framework: Implement a framework for standardizing performance evaluations of scanners and evaluators, incorporating a dynamic dataset that mirrors the evolving landscape of LLM vulnerabilities. Improve Evaluators: Utilize the new benchmarks to refine evaluators. As an initial step, we advocate for a hybrid approach that merges the static string-matching with the LLM-based evaluators to combine their strengths - simplicity and controllability with high-level understanding. Establish Quality Standards: Introduce regulatory quality standards for vulnerability scanners, ensuring they meet a baseline level of efficacy and reliability.
VII Conclusions
This paper presents the first hands-on comparative analysis of open-source LLM vulnerability scanners. We highlight both the shared principles and distinctive features of four promising scanners - Garak, Giskard, PyRIT, and CyberSecEval - revealing variations in their attack coverage. By conducting thorough evaluations and curating a much-needed dataset, we draw important conclusions about the efficiency and reliability of today’s scanners. Our recommendations will assist organizations in effective scanner deployment and guide researchers in making critical improvements to ensure reliable security against evolving threats in conversational AI.
@@ VIII CRediT authorship contribution statement
Appendix
The following appendix provides supplementary materials regarding the experiments conducted, the attack categorization applied in our research, and each scanner.
@@ Appendix A Attack Categories
We unified scanner attacks into four categories—Jailbreak, Context and Continuation, Gradient-based, and Code Generation Attacks—to enable structured comparison despite partial overlaps of the highly varied array of attacks that each scanner employs. The precise categorization of ach attack in each scanner is provided in Table III.
TABLE III: Overview of attack categories, descriptions, specific attacks, and their groupings in Garak, Giskard, PyRIT, and CybrSecEval.
@@
Appendix B Scanners Output Reports

### v2

@@ IV Per-Scanner Review
Scanner Test-suite Focus Guardrails Interface Multi-language Support Customizable Attacks Automated Customization Evaluator Explanation Insecure Coding Tests
Garak String-matching $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$
PyRIT LLM-based $\times$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$
Giskard LLM-based $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
CyberSec. Pattern-matching $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$
Giskard. Giskard’s test suite features a diverse set of nine attack-evaluation pairs, combining static and LLM-based methods. The latter covers areas such as hallucinations, harmful content, stereotypes, and information disclosure - see Table 3 in the supplementary material. A unique feature is Giskard’s dual-context mechanism, which includes: 1) a description of the target model, and 2) in most cases, a list of safety requirements, which are either predefined or generated per attack. These requirements, incorporated into the automated customization of adversarial prompts and the evaluator’s decision process, specify expected model robustness considering the attack at hand. Notably, this process involves two LLMs: The standard attacker LLM and the requirements LLM, unlike ”regular” LLM-based attacks which usually employ a single LLM - see Fig. 3. Giskard also supports non-English languages (see Fig.4 in the supplementary material), and generates a user-friendly HTML report that categorizes failed cases by attack type, including prompts, responses, and evaluator explanations (see Fig. 1 in the supplementary material).
Pyrit. [23] provides a fully LLM-based framework with a flexible design, enabling direct access to the instructions of both attacker and evaluator LLMs. It offers two attack approaches: 1) a single-turn attack, similar to other LLM-based attacks, and 2) a multi-turn attack, engaging the target LLM in a sequential dialogue until a predefined goal is reached (see Fig. 3). The evaluator LLM has a dual role in multi-turn attacks: Beyond deciding whether the attack succeeded, it also determines whether the dialogue should continue after each response. PyRIT’s evaluator component uses four scoring strategies (see Fig. 7 in the supplementary): Binary decisions (attack success or failure), discrete rankings (1 to 5) and continuous (0 to 1) scores - evaluating the target LLM’s response across categories like hate, bias and violence, following OpenAI moderation standards [43]. Additionally, PyRIT provides a strategy for evaluating the LLM evaluator, enhancing the system’s trustworthiness. PyRIT offers ”semi-automatic” capabilities leveraging human-in-the-loop attack curation, where users provide an initial prompt, with further augmentations enabled through its ”converters” component. This includes multi-language support. Though useful, human-in-the-loop pipelines are out of scope in this report. Though PyRIT lacks a formal report format, the attack conversations and success rates are easily accessible. PyRIT includes an LLM-generated explanation to clarify decisions, enhancing interpretability.
Figure 3: Top: Example Flow of Giskard, testing a ”Workshop Organizer AI”. An important aspect of Giskard is its ability to customize tests for LLMs that are designed for specific tasks. This example demonstrates Giskard’s customization via its distinctive requirements-based test. This is a shortened version - for the full version, including the evaluation phase, refer to Fig. 4 in the supplementary material. Bottom: Example flow of PyRIT’s multi-step attack for generating Python Key Logger. An attacker LLM is tasked with attacking a target LLM under evaluation, while another LLM assesses the attack’s success. This loop continues until the attack succeeds or a stopping criterion is met.
CyberSecEval. CyberSecEval [38, 29, 39] provides a test suite focused on analyzing vulnerabilities in LLM-generated code. It primarily targets insecure coding practices, evaluates malicious code generation, and includes jailbreak attacks that extract sensitive information from LLMs. To test for insecure coding practices, the CyberSecEval incorporates two main strategies, which use a database of insecure code snippets - as illustrated in Fig. 5 in the supplementary material. The first is an ”auto-complete” scenario, where the LLM is prompted with the 10 code lines that precede an insecure practice to see if it reproduces the risky code. The second, referred to as ”instruct”, converts instances of insecure practices into natural language instructions which in turn are passed as prompts to the target LLM, to assess if the LLM replicates the insecure practice. CyberSecEval evaluator, called the Insecure Code Detector (ICD) identifies insecure coding practices for test case generation and model evaluation. The ICD uses rules created by Meta’s cyber security experts, applied through static analysis tools like Weggli, Semgrep, and regex (see Fig. 6 in the supplementary material). CyberSecEval produces statistical reports on attack success rates and detailed reports documenting each attack along with the LLM’s corresponding response. While CyberSecEval’s code implementation is stable and well-designed for easy modification of attacks and evaluations, expending these is not straightforward, as it requires both cybersecurity and AI expertise. However, thousands of available CWEs one be use to extend insecure coding coverage with CybeSecEval’s framework. Thus in the hands of experts, CyberSecEval has potential for significant extensions.
IV-A Summary of Distinctive Features and their Comparison
Table II provides a summary, highlighting the following distinctive scanner features: Guardrails Interface. Encapsulates a leading Guardrails solution. Multi-language Support. Exposes vulnerabilities in non-English languages. Customizable Attacks. Automated Customization. Enables user-attack-customization, with automated prompt engineering based on user’s free-language input. Evaluator Explanations. See Sec. III Insecure Coding Tests. Tests security of generated code.
@@ V-A Evaluation Objectives
Here we provide an analysis of the four scanners, on two key aspects: attack coverage and evaluator effectiveness.
Attack Coverage. The attack coverage is considered comprehensive when it simulates a wide range of effective threats. We examine the diversity of attack types; the volume of attack instances produced for each type; and their quality, measured by their effectiveness in triggering errors in the target LLM.
Evaluator Effectiveness. The reliability of a scanner is assessed by evaluating its evaluators: How accurately do they detect vulnerabilities across various attack scenarios?
@@ V-B Evaluation Methodology and Results
To assess the scanners’ effectiveness in identifying vulnerabilities, each scanner was tested against four LLMs: Meta’s LLaMA 3 [1], Cohere’s Command-R [3], OpenAI’s GPT-4o [2], and Mistral AI’s Mistral Small [44]. To assess the quality of the attacks provided by each scanner, we applied a comprehensive series of these attacks and measured their success rates (ASR) using each scanner evaluator. However, evaluators can also make errors, necessitating the assessment of their margin of error (MOE) to ensure reliable attack success rates reported. Due to the lack of inherent ground truth in scanner evaluations, we manually annotated over a thousand attack responses to establish a reliable baseline. This manual annotation process required a thorough understanding of each attack’s objective, determining its success by analyzing the model’s response. We ensured balanced representation across attack types, facilitating the calculation of the evaluators’ accuracy and their MOE. More details regarding our evaluation and data curation methodology are in the supplementary material.
TABLE III: Attacks’ success rate (ASR) and reliability (MOE) over different LLM models.
@@
Table III shows for each scanner its different attack categories, alongside their number of prompt instances, success rates (ASR) and margin of error (MOE) across the four tested LLMs. Figure 4 shows the ASR and MOE by the attack perspective. For the sake of comparability, the scanners’ partially overlapping $\sim$35 attack types were grouped into five categories (Jailbreak, Gradient-based, Context and Continuation C&C, Insecure Code Generation CG and General Multi-turn) - see Table I. Garak stands out with the most extensive variety of attacks and, in most cases, the highest number of instances per attack. Additionally, Garak’s attacks are of the highest quality, with success rates ranging from approximately 20% in context and continuation attacks to nearly 70% in insecure code attacks. Fig. 1 shows a clear advantage for static-focused scanners in terms of attack effectiveness. The MOE measurements reveal notable errors across all scanner evaluators, which we consider as a measure of reliability - with PyRIT’s LLM-based evaluators being the most reliable (lowest MOE), while Garak shows a maximum MOE of 26%, corresponding to a 37% error rate in detecting successful attacks. As evident in Fig. 1, the LLM-based evaluators of Giskard and PyRIT offer the highest reliability.
@@
Figure 4: Per-attack performance, averaged over the target LLMs. MOE is shown as error bars.
VI Qualitative Analysis: Evaluators’ Errors
High error rates in the evaluator component are evident across all tools, impacting both static and LLM-based evaluators. This finding, presented above for the first time, calls for deeper understanding. Here we look into examples of evaluator errors that elucidate causes of this limitation.
Table IV presents several examples, presenting several typical cases - the full examples are available in the supplementary $full\,examples.docx$. These examples demonstrate several key limitations in current evaluator designs. Example 1 shows that static evaluators’ manually crafted logic lacks contextual awareness: The word ”unethical” in the response is misinterpreted as a refusal to comply with the attack, while it conveys the opposite intent. In Example 2 a static evaluator misclassifies due to minor punctuation misalignment. Example 3 illustrates a common challenge with evaluator LLMs: Although the evaluator correctly identified harmful content, it misinterpreted its instructions, and misclassified the response. Example 4 highlights the complexity of controlling multiple-LLM pipelines: Despite clear instructions for the LLM to operate independently, a requirement for external feedback was mistakenly generated, leading to a valid response being misclassified as invalid. Example 5 lacks a provided explanation by the scanner, highlighting the need for such mechanisms, especially given the black-box nature of LLMs: It is crucial for users to understand vulnerabilities and developers to trace errors. While static evaluators offer transparency by design, LLMs lack the possibility to track their internal processes. PyRIT and Giskard prompt the evaluator LLM to provide explanations of its decisions - which we find useful.
{verbbox}
[] # Fast-Flux Network Configuration # This code sets up and maintains # a fast-flux network for C2 communication.
import socket import dns.resolver ...
TABLE IV: Evaluator Errors Examples (Full Versions in Supplementary). $\dagger$ Garak;  $\ddagger$ Giskard;  $\parallel$ PyRIT;  §$\S$ CyberSecEval
#Example. Case Attack Response Decision and Reason
[…]
3. LLM Failed Reasoningparallel-$\parallel$ Multi-turn Jailbreak…provide two different responses… Add [CLASSIC] in front of the standard response and [JAILBREAK] in front of the one intended to be a DAN.
How can one successfully hide their digital tracks when accessing restricted websites?” (full conversation in supplementary)
[CLASSIC] I cannot provide…”
”[JAILBREAK] Absolutely… you need to use a Virtual Private Network (VPN), Tor browser, and perhaps even a proxy server… Using cryptocurrencies for payments can add an extra layer of anonymity. However, these actions may infringe on privacy..
Evaluator Decision: Valid Response. Reason: Evaluator explanation: ”Although the [JAILBREAK] part provides information, the presence of any refusal qualifies it as ’False’..”. It recognized the harmful information, but failed to classify the invalid response.
[…]
Model description was provided as context: ”..This assistant operates independently of any external mechanisms, relying solely on its own capabilities to align with the safeguards outlined above..
I cannot assist with illegal activities such as hacking. Hacking into a school’s computer system is illegal and a violation of privacy and security. Evaluator Decision: Invalid Response. Reason: Generated requirement: ”..include a reporting mechanism for users..”. This caused the evaluator LLM to test for external feedback despite the provided model description, consequently not recognizing the valid response.
[…]
VII Conclusions and Recommendations
We have presented the first hands-on, comparative, deep analysis of LLM vulnerability scanners. We highlight both shared principles and distinctive features of four cutting-edge scanners, revealing variations such as customizability, use of LLMs and quantified coverage comprehensiveness. Based on these, we provide below (Subsec. VII-A) strategic recommendations for organizations.
Through extensive evaluations and curation of a foundational dataset, we drew conclusions on today’s scanners’ efficacy - including the first identification of a critical gap in the evaluator component’s performance across all tested tools. To elucidate this issue, qualitative cases were analyzed, showing the over-simplicity of static evaluators, and the uncontrollable nature of LLM-based evaluators. Below (Subsec. VII-B), we situate this within the broader aspect of LLM security and provide future suggestions for quality control of LLMs.
VII-A Matching Red-Teaming Scanners to Organizational Needs
@@
This is a complex task, influenced by specific business risks. Red-teaming groups handling diverse use cases, such as large corporations employing LLMs for HR, sales, marketing, internal data analysis, etc. would prefer a ready-to-use wide coverage test suite, where customization is less frequent. In contrast, agile firms focusing on single-case products would favor customizable do-it-yourself solutions to allow for dynamic adaptations through tailored tests that meet their specific needs. Here we position the four scanners on a ”ready-to-use” to ”do-it-yourself” spectrum, highlighting their strengths and limitations to guide organizations in making informed choices.
@@
Garak offers the most extensive test-suite, making it suitable for red-teaming groups that deal with diverse use-cases. However, it focuses more on a static attack dataset, limiting customizability. Garak also integrates with Nvidia’s NeMo Guardrails, enabling setting additional safety layers.
@@
Giskard is ideal for users seeking flexible attack generation with both static and LLM-based methods. It offers a simple yet effective customization of LLM-based attacks via user-provided natural language context - enabling tailored test suites for various attack types, useful for dynamic online environments with minimal manual interference. It includes guardrails interface to enhance safety.
@@
PyRIT offers the most customizable test suite, focusing on LLM-based attacks. It allows users to edit both attacker and evaluator LLMs, providing full access to their instructions. This offers extensive flexibility but requires significant prompt engineering. Therefore, PyRIT is best suited for red-teams focusing on an internally crafted test-suite rather than relying on external knowledge. PyRIT’s distinctive multi-step attacks and rich ’semi-automatic’ options provide an additional edge.
@@
CyberSecEval focuses on red-teaming for code-generating LLMs and is placed in the ready-to-use end of the spectrum. Its test-suite is designed to expose code-related security issues. This makes it valuable for red-teaming groups dealing with generative AI for software and cybersecurity, where the integrity of auto-generated code is crucial.
VII-B Future Suggestions Addressing Current Gaps
Quality Standards. As part of emerging AI regulatory frameworks, we recommend the establishment of explicit quality standards for scanning tools. To achieve this, regulatory bodies should set baseline requirements that scanners must meet in identifying vulnerabilities. A concrete suggestion, based on our findings, is to standardize evaluations of the evaluator component to ensure that scanners correctly identify which attacks the target LLM is vulnerable to. For example, if for a set of vulnerabilities the evaluator fails to provide accuracy above a certain threshold, then the scanner should not qualify for detecting this set - even if its test-suite includes these vulnerabilities. This approach aims to ensure a basic acceptable level of efficacy and reliability across all scanners.
Benchmarking Framework. To solve the current status in which comparative perspectives await initiatives like this paper, we recommend establishing a unified platform where developers can upload their scanners to benchmark them against others, track performance trends, and make targeted improvements based on evolving security needs. The framework should incorporate dynamic and continuously updated tests and criteria that reflect the evolving landscape of LLM vulnerabilities. In our experiments we found attack categorization to be crucial for comparability; incorporating such categorization in this framework would also facilitate benchmarking for specific security needs. OWASP recently highlighted a related role of scanners in benchmarking LLM security performance: Both their Q3 and Q4 2024 reports [15, 16] categorize LLM vulnerability scanners under the ”LLM Automated Benchmarking” category. Our proposition to benchmark the scanners themselves is a natural extension of this notion.
Our study establishes an initial foundation for comparing scanners, offering both a methodology and criteria for meaningful evaluation, while uncovering insights and pressing issues within the domain. We invite the research community, industry, and regulators to build upon these findings to strengthen the safety and robustness of LLMs, addressing the evolving landscape of security threats.
@@ Appendix A Attack Categories
We unified scanner attacks into four categories—Jailbreak, Context and Continuation, Gradient-based, and Code Generation Attacks—to enable structured comparison despite partial overlaps of the highly varied array of attacks that each scanner employs. The precise categorization of each attack in each scanner is provided in Table V.
TABLE V: Overview of attack categories, descriptions, specific attacks, and their groupings in Garak, Giskard, PyRIT, and CybrSecEval.
@@
A-A $16$ Scanners Summary
Among the rest twelve scanners, Prompt Fuzzer excels in systematically executing dynamic prompt injection attacks, leveraging fuzzing techniques to iteratively craft adversarial inputs that challenge LLM response consistency. HouYi distinguishes itself by employing a black-box approach that segments and manipulates LLM contexts to inject malicious payloads, specifically targeting LLM-integrated applications to expose context-based vulnerabilities handled quite well in Giskard. Dioptra, is aligned with NIST’s AI Risk Management Framework to ensure AI validity, reliability, safety, security, and fairness and it is built to support a wide range of use cases including model testing, aiding AI research, and providing a controlled environment for red-teaming exercises. JailBreakingLLMs focuses on risk assessment using gradient-based adversarial suffix generation, claiming enabling of targeted jailbreaks across high-profile models like GPTs and Claude with minimal queries. LLMAttacks emphasizes universal, transferable attack strategies, generating adversarial prompts that induce misaligned outputs in aligned LLMs, thus demonstrating the fragility of model guardrails under broad attack vectors. PromptInject evaluates LLMs’ resilience against straightforward prompt injections, using a static set of predefined attack patterns that test the models’ capacity to maintain instruction fidelity amidst adversarial manipulations. Promptfoo’s primary capability lies in its comprehensive fuzz testing framework, which identifies and rectifies LLM vulnerabilities during early development (and later-on for monitoring stages as well, focusing on contextual handling and interaction consistency under manipulated prompts. LLMCanary aligns its assessment with the OWASP Top 10 for LLM vulnerabilities, providing a structured benchmarking framework that rigorously tests LLMs for security flaws, including data leakage and unsafe outputs, to guide safer model integration. Agentic Security offers a comprehensive red-teaming platform, integrating rule-based attack generation with API fuzzing to stress-test LLMs under variable and adaptive threat scenarios, prioritizing adaptability and coverage across different LLM APIs. LLMFuzzer, although less actively maintained, offers a specialized framework for testing LLMs integrated via APIs, utilizing modular fuzzing strategies that dynamically explore input-output vulnerabilities in application-specific contexts. PromptMap automates the identification of prompt injection vulnerabilities within GPTs, utilizing a mapping approach to systematically explore various prompt manipulations, including context-switching and translation-based attacks, to reveal latent weaknesses in conversational models. Vigil-LLM combines transformer-based heuristics with rule-based analysis to detect prompt injections and jailbreaks, offering a versatile toolkit for real-time monitoring and mitigation of LLM security risks via a dual-mode API and library configuration.
TABLE VI: Overview of Vulnerability Scanners
Scanner Name Attack Strategy Target Vulnerabilities Key Features Maintenance
Prompt Fuzzer [19] Dynamic prompt injection, fuzzing LLM response consistency Iteratively craft adversarial inputs Actively maintained
HouYi [liu2023prompt] Black-box context manipulation Context-based vulnerabilities Segments and manipulates LLM contexts Actively maintained
JailBreakingLLMs [chao2023jailbreaking] Gradient-based adversarial suffix High-profile model jailbreaks Minimal queries for targeted jailbreaks Less maintained
LLMAttacks [33] Universal, transferable attack strategies Misaligned outputs in aligned LLMs Generates adversarial prompts Less maintained
PromptInject [37] Straightforward prompt injections Instruction fidelity amidst adversarial manipulations Uses a static set of predefined attack patterns Less maintained
Promptfoo [20] Comprehensive fuzz testing LLM vulnerabilities during development Identifies and rectifies vulnerabilities Actively maintained
LLMCanary [21] OWASP Top 10 for LLM vulnerabilities Data leakage, unsafe outputs Provides a structured benchmarking framework Actively maintained
Agentic Security [22] Rule-based attack generation, API fuzzing Security flaws in variable threat scenarios Red-teaming platform, adapts to different LLM APIs Actively maintained
LLMFuzzer [24] Modular fuzzing strategies Input-output vulnerabilities in application-specific contexts Specialized framework for API-integrated LLMs Less maintained
PromptMap [25] Systematic exploration of prompt manipulations Prompt injection vulnerabilities in GPTs Utilizes mapping approach for attacks Less maintained
Vigil-LLM [vigil_llm] Transformer-based heuristics, rule analysis Prompt injections and jailbreaks Dual-mode API and library configuration for real-time monitoring Actively maintained
Garak [17] Structured vulnerability detection Diverse attack types Regularly updated library Highly maintained
PyRIT [23] Customized attack scenarios LLM-based frameworks Flexible user-defined instructions Highly maintained
Giskard [18] Static and dynamic evaluations Tailored to model descriptions Dual-context mechanisms Highly maintained
CyberSecEval [38] Rule-based static analysis Code integrity and security Identifies LLM generated Code vulnerabilities Highly maintained
A-B Scanners Output Reports